OpenStudy (snuggielad):

Showing someone how the site works 3|x-5|=20 Solve For X Please give me some assistance @amistre64 @thomaster

OpenStudy (amistre64):

divide; flip out on negatives, and add

OpenStudy (amistre64):

\[m|x+c|=P\] \[|x+c|=\frac Pm\] \[-\frac Pm=(x+c)=\frac Pm\] \[-\frac Pm-c=(x)=\frac Pm-c\]

OpenStudy (snuggielad):

Thanks! @Ryan_Kochan

OpenStudy (snuggielad):

Well, thanks @amistre64 but I am tagging @Ryan_Kochan to let him see...

OpenStudy (amistre64):

notation might be better if we used:\(\leftarrow(x)\rightarrow\)